OpenStudy (anonymous):

If k is an interger and numbers 1+k, 1+2k and 1+8k are divisible by 3,5 and 7 respectively, then the smallest positive value of k lies in the interval a) 0 to 20 b) 21 to 40 c) 41 to 60 d) 61 to 80 e) none of these

OpenStudy (asnaseer):

I think you can use the Chinese Remainder Theorem here. See here for a good tutorial on this: http://www.youtube.com/watch?v=3PkxN_r9up8

OpenStudy (asnaseer):

I have worked out the answer. Let me know what answer you get and I'll tell you if it matches the one I found using this theorem.

OpenStudy (anonymous):

i think i get the chinese remainder theorem but i am not sure how to apply it in this problem

OpenStudy (asnaseer):

ok, lets take each part that you have been given. 1+k is divisible by 3 means:\[1+k=0\text{ mod }3\]\[\therefore k=-1\text{ mod }3=2\text{ mod }3\]similarly 1+2k is divisible by 5 means:\[1+2k=0\text{ mod }5\]\[\therefore 2k=-1\text{ mod }5=4\text{ mod }5\]\[\therefore k=2\text{ mod }5\]and lastly, 1+8k is divisible by 7 means:\[1+8k=0\text{ mod }7\]\[\therefore 8k=-1\text{ mod }7=6\text{ mod }7\]\[\therefore 4k=3\text{ mod }7\]\[\therefore -3k=3\text{ mod }7\]\[\therefore l=-1\text{ mod }7=6\text{ mod }7\] so we end up with:\[k=2\text{ mod }3\]\[k=2\text{ mod }5\]\[k=6\text{ mod }7\]Now just use the theorem to find the general solution for k.

OpenStudy (anonymous):

i got 452 for k

OpenStudy (asnaseer):

you need to get the general solution, something like: k = a + nb where a and b are integers and 'n' is a variable multiplier that can range from \(-\infty\) to \(+\infty\). you can then use that formula to work out the minimum positive value for k.

OpenStudy (asnaseer):

452 is not one of the valid values for k.

OpenStudy (anonymous):

k = 260 +/- n105

OpenStudy (asnaseer):

the n105 is correct but the 260 is wrong.

OpenStudy (asnaseer):

I need to leave for a wedding now - but if you write out the steps you took here then when I get back I can check them to see where you made a mistake.

OpenStudy (anonymous):

o i find my mistake now i get k = 272 +/- 105 and the smallest value is 62

OpenStudy (anonymous):

thanks asnaseer for that info - interesting stuff

OpenStudy (asnaseer):

yociyoci - that is correct - well done for pursuing this. jimmyrep - yw

OpenStudy (anonymous):

Thank you very much for showing this new theorem to me and showing me the steps to this problem. Thank you!!!!

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

asnaseer, i have a problem posted earlier, and it was left unfinish, can you help me out? Thanks. http://openstudy.com/study#/updates/4f3bcab2e4b0fc0c1a0f1d10

OpenStudy (asnaseer):

I haven't seen a problem like that one before - let me think over it - I'll get back to you.

OpenStudy (anonymous):

okay, thank you.